OpenStudy (gavin39):

Will give medal and become friend if you could please help me answer this Solve 3x−5 = 9. x = −3 x = 2 x = 7 x = 8

OpenStudy (anonymous):

the goal is to get x alone while remembering that what you do to one side you have to do to the other!

OpenStudy (anonymous):

can you start by removing the 5 from the side that has the x? show me what you got!

OpenStudy (anonymous):

but... actually, none of those answers are correct. did you type in the question properly?

OpenStudy (gavin39):

Yeah straight from the question online (im virtual school). We are left with 3=4

OpenStudy (gavin39):

That was I was thinking it didn't work right

OpenStudy (anonymous):

we get 3x-5 +5 = 9 +5

OpenStudy (gavin39):

the only one I can think of it being is 2

OpenStudy (anonymous):

3x - 0 = 14 are you able to see why?

OpenStudy (gavin39):

wait are you using pemdas?

OpenStudy (gavin39):

O wait you sitched the 5 and 9 I see

OpenStudy (anonymous):

i am moving stuff around so that x can be alone on 1 side. 3x-5 = 9 right? yes!

OpenStudy (gavin39):

ah ok I see!

OpenStudy (anonymous):

once you have 3x = 14 how could you get rid of that 3 on the left side?

OpenStudy (anonymous):

14/3 is still not any of those

OpenStudy (anonymous):

i know, im trying to just show him how to solve these, im thinking he left off a sign somewhere.